OpenStudy (anonymous):

Graph each pair of parametric equations 1.x=3 sin^3 t y=3 cos^3 t

OpenStudy (anonymous):

\[\frac{ x }{3 }=\sin ^{3}t,\] taking cuberoot we get \[\left( \frac{ x }{ 3 } \right)^{\frac{ 1 }{3 }}=\sin t\] squaring both sides \[\left( \frac{ x }{3 } \right)^{\frac{ 2 }{3 }}=\sin ^{2}t ...(1)\] similarly \[\left( \frac{ y }{ 3 } \right)^{\frac{ 2 }{3 }}=\cos ^{2}t ...(2)\] adding (1) & (2) \[\left( \frac{ x }{3 } \right)^{\frac{ 2 }{3 }}+\left( \frac{ y }{3 } \right)^{\frac{ 2 }{ 3 }}=\sin ^{2}t+\cos ^{2}t=1\] now you can make the graph.